OpenStudy (anonymous):

How would you set up an equation for this? A party-length submarine sandwich that is 59 in. long is cut into three pieces. The middle piece is 5 in. longer than the shortest piece, and the shortest piece is 9 in. shorter than the longest piece. How long is each piece?

OpenStudy (anonymous):

Well in the equation, just solve for x: x + (9+5) + 9 = 59 x + (14) +9 = 59 x + 23 = 59 x = 59 - 23 x = 36 first part: 36 middle: 14 last part: 9 total: 59

OpenStudy (anonymous):

I don't think that answer is right

OpenStudy (anonymous):

Let's call the shortest piece is x then the middle piece is x + 5 and the longest piece is x +9 add them all up equals to 59 to get the equation

OpenStudy (anonymous):

So the equation would be x+x+5+x+9=59?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

combine like-terms now

OpenStudy (anonymous):

oh ok, that makes it easier. I had x as the longest piece, I thinks thats were i want wrong... Thank you!!!

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

I apologize I read it wrong